# Exploring the interaction between the MW and LMC with a large sample of blue horizontal branch stars from the DESI survey

### v1

@@ 4 Model comparison
When making our comparisons between these models and the data, we apply the following on-sky selection to the models to mimic the DESI footprint:
$|b|>20^{\circ}$ (11)
and
The DESI footprint and the X11 footprint, with the latter being SDSS data, are almost identical, and so we assume that the DESI footprint is representative of the entire combined BHB sample.
We first wish to understand if the BHB $v_{\text{GSR}}$ distribution as a function of distance behaves similarly to these models. This is done by computing the mean $v_{\text{GSR}}$ and its error in 10 kpc wide Galactocentric radial bins, using Eqs. (8) and (9), for the Northern and Southern hemispheres separately, for both the BHB sample and the halo stellar particles in all of the Vasiliev (2024) models. The comparison between the BHB sample and the models is shown in Fig. 10. In the figure, the mean $v_{\text{GSR}}$ for the BHB sample is shown as connected points with corresponding error bars in red and blue for the Northern and Southern hemispheres respectively. The same for the models are shown as dashed lines. Looking at this figure, we see that all models have the same qualitative behaviour in velocity as a function of distance, and they all fail to properly fit the data. In Table 2, we compare the reduced $\chi^{2}$ of each model to the data by comparing the bins in Fig. 10. This comparison shows that the L2M11 model is the best fit, and for this reason alone, we will use it for comparison to our data throughout the rest of the paper.
@@
If we again look at all the stars in the best-fitting model L2M11 that have Galactocentric distances between 50 and 120 kpc (where the upper range reflects the distance distribution of our sample, see Fig. 4), and use the same Voronoi bins as in Fig. 6, we can construct its on-sky velocity distribution. This is seen in Fig. 12, where the bins that are empty in the version for the BHB are also empty for the model because we apply the footprint cut in Eq. (11).
@@
To understand if the discrepancy between the model and the data really is due to differences in compression velocity, we apply the velocity decomposition of Eq. (10) to the L2M11 model, over the distance range 0–120 kpc and with the on-sky selection given in Eq. (11) applied. Doing so gives us the values in Table 3, which comparing to the corresponding values for the BHB sample shown in Table 1 shows that the data gives a $v_{\text{compr}}$ that is $\sim 10$  km s-1 larger in amplitude and a $v_{\text{dipole}}$ that is $\sim 10$  km s-1 smaller in amplitude than the model. This means that the model’s velocity field more purely consists of a dipole than the data’s velocity field. This explains why the model and the data agree in the Southern hemisphere in Figs. 11 and 10, where the two velocity components are added, and the $\sim 20$  km s-1 difference between model and data in the North. We also note that the apex direction for L2M11 agrees much better with our work than it does for PP21, Y24. The angular offset in apex direction compared to our results differ by about 32, and it differs from the directions by PP21 and Y24 by 47 and 45 respectively. This angular offset to PP21 was pointed out already by Vasiliev (2024, their Sec. 3.2), when comparing the sky-projected velocity fields.
@@
$v_{\text{compr}}$ [ km s-1] $-10.1\pm 0.1$
$v_{\text{dipole}}$ [ km s-1] $-39.1\pm 0.2$
$l_{\text{apex}}$ [deg] $-18.5\pm 2.3$
$b_{\text{apex}}$ [deg] $-80.8\pm 0.6$
$\sigma$ [ km s-1] $98.8\pm 0.1$
@@ 5.2 Understanding the compression velocity
Our parametrization of the velocity field does not contain a local wake component, and if it is present in the data but not in the model, the expected velocity field from Eq. (10) subtracted from the data should reveal a negative spot at the location of the local wake. We test this by redoing the fit of Eq. (10) on all models by Vasiliev (2024), and subtract this fitted velocity field from the simulated velocity data. This is shown for the L2M11 model in Fig. 10, that also includes the DESI footprint as given by Eq. (11). This reveals a region with velocity more negative than the dipole prediction from our model, centered at $l\approx 75^{\circ}$ for all six models, with a center in $b$ varying between $\sim-10^{\circ}$$0^{\circ}$. The amplitude of this region is $\sim 10-20$  km s-1. This range roughly encompasses our measured velocity compression of –23.5  km s-1. We do not see particularly strong negative residuals in our model in the $(l,b)\approx(75^{\circ},-15^{\circ}$) region, see Fig. 15, suggesting that the potential presence of the local wake does not affect our data nor the model fit to it.
@@ 5.3 Comparison with literature apex directions
The BHB sample presented in this work is the largest spectroscopic BHB sample in the literature, and will only grow with upcoming DESI data releases. The DESI BHB sample contains 9,866 BHBs, and SDSS DR6 contains 2,401 BHBs (Xue et al., 2008) and SDSS DR8 contains 4,985 BHB stars (X11). The catalogue’s strengths is not only its increased number of BHB stars compared to previous spectroscopic works: the DESI survey also reach fainter magnitudes leading to a larger probed distance range. It also has smaller velocity errors than previous works. That we only use BHB stars instead of combining different halo tracers, such as RGB stars, leads to a less complex sample, that is homogeneous across the entire probed distance range. Comparing the apex direction we derive using our BHB sample to that published in Chandra et al. (2024), we are in agreement. They use K-giants across the entire sky and we instead have limited sky coverage (see Fig. 5). However, if we compare our direction to those published by PP21 and Y24 using spectroscopic BHB tracers from SDSS, our dipole directions disagree, with an angular offset of $76^{\circ}$ and $57^{\circ}$ respectively (as seen in Fig. 9). Interestingly, the amplitude of the velocity fields agree. Because of this, there is rough agreement in predicted halo velocities where $l>0^{\circ}$, but not complete agreement across the celestial sphere, and this lack of complete agreement is due to the differences in apex directions. We must understand where this disagreement comes from: is it due to our different data sets, or because of the different models we use for our velocity fields?
@@
The other possible explanation for the tension of results is differences in the BHB samples. The works by PP21 and Y24 use K-giant and BHB tracer samples, but we continue to focus on their results derived using BHBs. Their BHB samples are based on the SDSS sample by X11, which is the same sample we use for the SDSS portion of our sample. Comparing their BHB sample to the combined one used in this work, we see that their sample is different from ours in distance and spatial distribution for stars with Galactocentric distance larger than 50 kpc. We can see in Fig. 4 that their sample is more concentrated around the Galactic centre and that we have more stars at greater distances: they have 2.2  per cent of stars past 50 kpc (7.1  per cent past 40 kpc), while we have 7.8  per cent past 50 kpc (or 16.1  per cent past 40 kpc). The two samples have roughly the same sky coverage, but their work measures more stars with $l<-100^{\circ}$, which is the main region in which our predicted velocities differ (and they mostly agree where $l>0^{\circ}$), and our work is measuring more stars in the Southern hemisphere ($b<0^{\circ}$). The differences in distance and spatial distributions between the data sets are thus subtle nd difficult to isolate, but we can see from the appendix in Y24 that that could still affect the sample.
@@
The final major difference between our data set and those of PP21, Y24 is the masking of Sgr stars. We compare the two methods of selecting Sgr stars on the DESI BHB stars in Appendix D. The comparison shows that the method used by PP21, Y24 leaves Sgr stars in our BHB sample as visible overdensities, especially at the Sgr leading arm apocentre, and all of the Cetus-Palca stream (we note that Cetus-Palca is not present in their data and that it is left in our sample does not imply that it is also left in theirs). The stream stars that are not removed from our BHB sample and at Galactocentric distances larger than 50 kpc, and thus enter into the analysis, have largely positive radial velocities and lie in the Northern hemisphere, see Fig. 5 where the Sgr stream track is shown in purple. The treatment of Sgr stars in our samples thus appears to be a major source of differences in our apex directions, and the way it is removed from the working sample will systematically bias the point of the apex direction due to the location of the stream in the Northern hemisphere.
The apex directions measured in this work and in Chandra et al. (2024) are in excellent agreement, see Fig. 9. We believe that a major discrepancy between our work and that by PP21, Y24 is due to our differences in selecting Sgr stream stars (see Appendix D) where they use cuts based on angular momentum. However Chandra et al. (2024) also use a cut in angular momentum but instead of selecting a sphere in $(L_{x},L_{y},L_{z})$ like PP21, Y24, they use the linear selection in $(L_{y},L_{z})$ proposed by Johnson et al. (2020). They supplement the angular momentum cuts with cuts in Sgr latitude $\tilde{B}$ and line-of-sight distance to the stream, which is similar to what we do in Eq. (7), as well as a velocity cut. This approach is similar enough to ours to not cause discrepancies, and applying this removal to our data still returns an apex direction that is in agreement with the one reported in this paper.
@@ 6 Summary and conclusions
  • Our apex direction is in good agreement with that by Chandra et al. (2024) but significantly far away from measurements by PP21 and Y24. We believe that the latter disagreement is driven by different methods in removing the Sgr stream stars from the outer halo samples, with the method presented in this work leading to a purer smooth halo sample.
@@ Acknowledgements
The data behind all figures in this paper will be made available as .fits files upon acceptance of the paper.

### v3

@@ 4 Model comparison
When making our comparisons between these models and the data, we apply the DESI footprint as seen in grey in Fig. 5 to the models to mimic the DESI footprint [^13] A tutorial on how to do this footprint selection as well as the DESI footprint file can be found at: https://github.com/abystrom/apex-likelihood-fitting.. The DESI footprint and the X11 footprint, with the latter being SDSS data, are almost identical, and so we assume that the DESI footprint is representative of the entire combined BHB sample.
We first wish to understand if the BHB $v_{\text{GSR}}$ distribution as a function of distance behaves similarly to these models. This is done by computing the mean $v_{\text{GSR}}$ and its error in 10 kpc wide Galactocentric radial bins, using Eqs. (8) and (9), for the Northern and Southern hemispheres separately, for both the BHB sample and the halo stellar particles in all of the Vasiliev (2024) models. The comparison between the BHB sample and the models is shown in Fig. 10. For Galactocentric distances below 20 kpc, the halo rotation significantly biases fits of Eq. (10), as shown in Appendix E. This means that the halo rotation is large at those distances. For this reason we only show Galactocentric distances above 20 kpc in the figure, even though the BHB sample extends down to 6 kpc. In the figure, the mean $v_{\text{GSR}}$ for the BHB sample is shown as connected points with corresponding error bars in red and blue for the Northern and Southern hemispheres respectively. The same for the models are shown as dashed lines. Looking at this figure, we see that all models have the same qualitative behaviour in velocity as a function of distance, and they all fail to properly fit the data. In Table 2, we compare the reduced $\chi^{2}$ of each model to the data by comparing the bins in Fig. 10, for Galactocentric distances above 50 kpc. This comparison shows that the L2M11 model is the best fit, and for this reason alone, we will use it for comparison to our data throughout the rest of the paper.
@@
If we again look at all the stars in the best-fitting model L2M11 that have Galactocentric distances between 50 and 120 kpc (where the upper range reflects the distance distribution of our sample, see Fig. 4), and use the same Voronoi bins as in Fig. 6, we can construct its on-sky velocity distribution. This is seen in Fig. 12, where the bins that are empty in the version for the BHB are also empty for the model because we apply the DESI footprint.
@@
To understand if the discrepancy between the model and the data really is due to differences in compression velocity, we apply the velocity decomposition of Eq. (10) to the L2M11 model, over the distance range 0–120 kpc and with the DESI footprint applied. Doing so gives us the values in Table 3, which comparing to the corresponding values for the BHB sample shown in Table 1 shows that the data gives a $v_{\text{compr}}$ that is $\sim 10$  km s-1 larger in amplitude and a $v_{\text{dipole}}$ that is $\sim 10$  km s-1 smaller in amplitude than the model. This means that the model’s velocity field more purely consists of a dipole than the data’s velocity field. This explains why the model and the data agree in the Southern hemisphere in Figs. 11 and 10, where the two velocity components are added, and the $\sim 20$  km s-1 difference between model and data in the North. We also note that the apex direction for L2M11 agrees much better with our work than it does for PP21, Y24. The angular offset in apex direction compared to our results differ by about 30, and it differs from the directions by PP21 and Y24 by 47 and 45 respectively. This angular offset to PP21 was pointed out already by Vasiliev (2024, their Sec. 3.2), when comparing the sky-projected velocity fields.
@@
$v_{\text{compr}}$ [ km s-1] $-11.2$
$v_{\text{dipole}}$ [ km s-1] $-40.5$
$l_{\text{apex}}$ [deg] $-25.3$
$b_{\text{apex}}$ [deg] $-78.9$
$\sigma$ [ km s-1] $98.3$
@@ 5.2 Understanding the compression velocity
Our parametrization of the velocity field does not contain a local wake component, and if it is present in the data but not in the model, the expected velocity field from Eq. (10) subtracted from the data should reveal a negative spot at the location of the local wake. We test this by redoing the fit of Eq. (10) on all models by Vasiliev (2024), and subtract this fitted velocity field from the simulated velocity data. This is shown for the L2M11 model in Fig. 15, that also includes an approximation of the DESI footprint as given by:
$|b|>20^{\circ}$ (11)
and
This reveals a region with velocity more negative than the dipole prediction from our model, centered at $l\approx 75^{\circ}$ for all six models, with a center in $b$ varying between $\sim-10^{\circ}$$0^{\circ}$. The amplitude of this region is $\sim 10-20$  km s-1. This range roughly encompasses our measured velocity compression of –23.5  km s-1. We do not see particularly strong negative residuals in our model in the $(l,b)\approx(75^{\circ},-15^{\circ}$) region, see Fig. 15, suggesting that the potential presence of the local wake does not affect our data nor the model fit to it.
@@ 5.3 Comparison with literature apex directions
The BHB sample presented in this work is the largest spectroscopic BHB sample in the literature, and will only grow with upcoming DESI data releases. The DESI BHB sample contains 9,866 BHBs, and SDSS DR6 contains 2,401 BHBs (Xue et al., 2008) and SDSS DR8 contains 4,985 BHB stars (X11). The catalogue’s strengths is not only its increased number of BHB stars compared to previous spectroscopic works: the DESI survey also reach fainter magnitudes leading to a larger probed distance range. It also has smaller velocity errors than previous works. That we only use BHB stars instead of combining different halo tracers, such as RGB stars, leads to a less complex sample, that is homogeneous across the entire probed distance range. Comparing the apex direction we derive using our BHB sample to that published in Chandra et al. (2024), we are in agreement. They use K-giants across the entire sky and we instead have limited sky coverage (see Fig. 5). However, if we compare our direction to those published by PP21 and Y24 using spectroscopic BHB tracers from SDSS, our dipole directions disagree, with an angular offset of $76^{\circ}$ and $57^{\circ}$ respectively (as seen in Fig. 9). The aforementioned works utilise Gaia proper motions, while we solely use radial velocities. Interestingly, the amplitude of the velocity fields agree. Because of this, there is rough agreement in predicted halo velocities where $l>0^{\circ}$, but not complete agreement across the celestial sphere, and this lack of complete agreement is due to the differences in apex directions. We must understand where this disagreement comes from: is it due to our different data sets, or because of the different models we use for our velocity fields?
@@
The other possible explanation for the tension of results is differences in the BHB samples. The works by PP21 and Y24 use K-giant and BHB tracer samples, but we continue to focus on their results derived using BHBs. Their BHB samples are based on the SDSS sample by X11, which is the same sample we use for the SDSS portion of our sample. Comparing their BHB sample to the combined one used in this work, we see that their sample is different from ours in distance and spatial distribution for stars with Galactocentric distance larger than 50 kpc. We can see in Fig. 4 that their sample is more concentrated around the Galactic centre and that we have more stars at greater distances: they have 2.2  per cent of stars past 50 kpc (7.1  per cent past 40 kpc), while we have 7.8  per cent past 50 kpc (or 16.1  per cent past 40 kpc). PP21 used an outer halo sample with $R>40$ kpc, where we instead chose to use $R>50$ kpc. If we instead use $R>40$ kpc and redo the fit of the halo velocity field on the DESI BHB sample, our $(l_{\text{apex}},b_{\text{apex}})$ equals $(-65.2,-52.4)$ degrees, which is closer to the apex direction from PP21 than the one for $R>50$ kpc which we present in this paper, though only by about 2 degrees. The two samples have roughly the same sky coverage, but their work measures more stars with $l<-100$ degrees, which is the main region in which our predicted velocities differ (and they mostly agree where $l>0^{\circ}$), and our work is measuring more stars in the Southern hemisphere ($b<0^{\circ}$). If we subtract the predicted velocity fields from PP21 and Y24 from ours (Fig. 8), we also see that our data (Fig. 5) extends further into regions of larger discrepancy than their data does. The differences in distance and spatial distributions between the data sets are thus subtle and difficult to isolate, but we can see from the appendix in Y24 that that could still affect the sample.
@@
The final major difference between our data set and those of PP21, Y24 is the masking of Sgr stars. We compare the two methods of selecting Sgr stars on the DESI BHB stars in Appendix D. The comparison shows that the method used by PP21, Y24 leaves Sgr stars in our BHB sample as visible overdensities, especially at the Sgr leading arm apocentre, and all of the Cetus-Palca stream (we note that Cetus-Palca is not present in their data and that it is left in our sample does not imply that it is also left in theirs). The stream stars that are not removed from our BHB sample and at Galactocentric distances larger than 50 kpc, and thus enter into the analysis, have largely positive radial velocities and lie in the Northern hemisphere, see Fig. 5 where the Sgr stream track is shown in purple. The treatment of Sgr stars in our samples thus appears to be a significant source of differences in our apex directions, and the way it is removed from the working sample will systematically bias the point of the apex direction due to the location of the stream in the Northern hemisphere.
@@ 6 Summary and conclusions
  • Our apex direction is in good agreement with that by Chandra et al. (2024) but significantly far away from measurements by PP21 and Y24. We believe that the latter disagreement is to some degree explained by different methods in removing the Sgr stream stars from the outer halo samples.
@@ Acknowledgements
All DESI spectroscopically confirmed BHBs (see Sec. 2.4) and their distances (see Sec. 2.5) that are part of the DESI DR1 dataset are available at: https://data.desi.lbl.gov/doc/releases/dr1/vac/mws-bhb/. The DESI Y2 BHB subsample will be made available in DESI DR2. The data behind all figures in this paper are available as .fits files at: https://doi.org/10.5281/zenodo.13711675. The code used to fit the parameters in Eq. (10) and to choose a DESI-like sky footprint (see Sec. 4) are available at: https://github.com/abystrom/apex-likelihood-fitting/.
@@ Appendix D Compare Sgr selections
In Fig. 19, we see where the stars that are selected using this work’s method in Eq. (7) and the $L$ Sgr selection, and have $R>50$ kpc, are distributed in Galactic coordinates in purple (top and bottom panels, respectively). Behind these stars, in grey, is the rest of the BHB sample (so that the grey stars in the bottom panel are the same stars seen in Fig. 18). All of the Sgr-selected stars are in the Northern hemisphere (except for one using the $L$ selection method). Because the $L$ selection is incomplete and leave Sgr stars in the sample, behind the Sgr stars in the bottom panel, we can still see an overdensity. Most of these stars have positive velocities. Fitting the apex direction with these stream stars in the sample will rotate it towards a point further along the past orbit of the LMC, because the anti-apex direction in the Northern hemisphere, where the positive contributions from the dipole velocity peaks, will come closer to where these remnant Sgr stars are in the sample. This leads to an apex direction that is closer to those reported in PP21, Y24. If we take their data and apply their Sgr removal method to it, we recover their apex value. If we instead apply our Sgr removal method in Eq. (7) to their data, we recover an apex direction that is closer to ours.
@@
Appendix E Validating absence of halo rotation bias
The parametrization of the halo velocity field as given by Eq. (10) does not take the stellar halo’s rotation into account. A prograde halo rotation has been measured by Deason et al. (2017) as ranging from 5–25  km s-1 in a Galactocentric radial range from 0–50 kpc, by Dillamore et al. (2023) as $\sim 15$  km s-1 in the range 6–16 kpc, by Y24 as increasing from 11  km s-1 in a 20–30 kpc range to 24  km s-1 beyond 50 kpc, and by PP21 as being 20  km s-1 beyond 40 kpc. Chandra et al. (2024) measure a halo rotation that changes from retrograde to prograde as distance increases. It was shown by PP21 that a rotating halo will create a similar pattern projected on the celestial sphere using line-of-sight velocities as a pure dipole signal (see their Extended Data Fig. 8), and that using a limited observational footprint may make it difficult to disentangle the rotation signal from a dipole signal. They point out that the degeneracy between rotation and dipole signatures in velocity can be broken if transverse velocities are included in the fit, alongside radial velocities. Considering that the DESI footprint does not cover the entire sky (see Fig. 5) and that we only utilise radial velocities in this work, we need to verify that the presence of rotation in the MW halo does not introduce a bias in our apex direction and that the halo rotation is not absorbed into $v_{\mathrm{compr}}$ or $v_{\mathrm{dipole}}$.
We test the potential bias in our parameters as given by Eq. (10) by comparing the fit using the L2M11 model (which does not contain ab initio rotation) seen in Sec. 4, with a new fit after rotation has been projected onto the radial velocities along the line-of-sight. This is done first for all stellar particles past 50 kpc, and second we compare the fits in different distance bins, since we expect that the effect of rotation should increase as Galactocentric distances decrease. For the first test, we apply three different rotations: –20  km s-1 from PP21, –24  km s-1 from Y24, and 12  km s-1 at 40 kpc to –10  km s-1 at 120 kpc with a linear variation in between, from Chandra et al. (2024). The resulting apex directions are shown in Table 6, together with the angular separation between the case of no rotation (as seen in Sec. 4) and the different cases of added rotation, which we denote $\theta$. A larger value of $\theta$ would mean that the assumed rotation would have a larger effect on the derived apex direction, and the opposite is true for a smaller $\theta$. However, from the range of $\theta$ values we see that the projection of rotation onto the line-of-sight shifts the apex direction very little, and the angular differences are much smaller than the errors on $(l_{\text{apex}},b_{\text{apex}})$ from the DESI BHB sample, given in Table 1. For this reason, we believe that the rotation of the halo does not affect our velocity field measurement.
Table 6: The resulting $(l_{\text{apex}},b_{\text{apex}})$ from redoing the fits of Eq. (10) on the L2M11 by Vasiliev (2024), with varying amount of halo rotation projected onto the radial velocity, and the angular separation $\theta$ between the unmodified apex direction and those with added rotation.
Rotation source $(l_{\text{apex}},b_{\text{apex}})$ [deg] $\theta$ [deg]
No rotation $(-25.3,-78.9)$
PP21 $(-4.3,-79.9)$ 4.0
Y24 $(0.2,-79.9)$ 4.8
Chandra et al. (2024) $(-27.1,-78.8)$ 0.4
For the second test, we bin the L2M11 model in 5 kpc distance bins that range from 6 kpc to 120 kpc, to match the distance distribution of our BHB sample, and repeat the fit of Eq. (10) in each distance bin, for both the unmodified case and when the rotation from Y24 has been added to the line-of-sight. We choose their rotation measurement as their rotation amplitude in the outer halo is the highest, and it thus represents a worst-case scenario. This test should inform us of the magnitude of the bias introduced by rotation. Since their rotation is only measured down to a Galactocentric distance of 20 kpc, we supplement this with the rotation from Dillamore et al. (2023) in the range 6 to 20 kpc. The differences between the unmodified case and the rotation-added case can be seen in Fig. 20. These plots show that the effects of adding rotation indeed are the largest for smaller distance bins, and at $R<30$ kpc, the effects are substantial. This means that in the inner halo, we cannot accurately measure the dipole, as the results from fitting Eq. (10) will return biased results. However, beyond 50 kpc, where we perform our apex direction fits, the effects of rotation are completely negligible, and well within the measurement uncertainties of the fits in Table 1: the $v_{\mathrm{compr}}$ amplitude changes are less than 0.4  km s-1, the $v_{\mathrm{dipole}}$ amplitude changes are less than 2  km s-1, and the angular differences less than 10 degrees. At $R=59.9$ kpc, which is the median distance of our sample when $R>50$ kpc, the difference in $(l_{\text{apex}},b_{\text{apex}})$ is 6.6 degrees.
Figure 20: The absolute value of the difference in the compression velocity $v_{\text{compr}}$ (top row), dipole velocity $v_{\text{dipole}}$ (middle row) and the apex direction $(l_{\text{apex}},b_{\text{apex}})$ (bottom row) for the L2M11 model without rotation and rotation from Dillamore et al. (2023) in the Galactocentric radial range 6–20 kpc and from Y24 beyond 20 kpc. The distance range used, $1206<R<$ 120 kpc, is chosen to match that of the BHB distribution, see Fig. 4.
These tests show that since the bias from halo rotation being projected onto the line-of-sight and thus our radial velocities is significantly smaller than the measurement uncertainties of our fit of the halo velocity field, we assume that a rotating halo does not bias the results presented in this paper. The rotation bias could be even smaller than presented in the second test, which assumed the highest amplitude rotation in literature (Y24). It is also clear that the effects of rotation are significant in the inner parts of the MW, which confirms the importance of using distant halo tracers to understand the effects of the MW’s reflex motion on the halo velocity field.